# Kinematics of the Sagittarius Dwarf Spheroidal core: A 5D Analysis for a 6D Methodology with Gaia DR3

### v1

@@ Abstract
pubyear: 2025pagerange: Kinematics of the Sagittarius Dwarf Spheroidal core: A 5D Analysis for a 6D Methodology with Gaia DR3A
@@ 1 Introduction
The Sagittarius (Sgr) dwarf spheroidal (dSph) is one of the most luminous satellite galaxies orbiting the Milky Way (MW) (Ibata et al., 1994). It is unique from other dSphs in that it has both a central core and long, extended tails that indicate tidal interaction with the MW (Law & Majewski, 2010). The photometric and kinematic properties of the central core and tidal tails, combined with simulations of its orbit, provide insight into the nature of the Sgr progenitor before it was accreted into the MW, and provide estimates for its dark-to-luminous mass ratio (Peñarrubia et al., 2010; Łokas et al., 2010a). Sgr-like progenitors have also been identified in large-volume cosmological simulations (Łokas, 2024), which provide insight into its evolution within a cosmological context.
@@
Though the dark matter distributions of many dSphs have been examined in detail, there have been relatively few studies of Sgr, particularly in the context of equilibrium models (Łokas et al., 2010b). This is mainly because it is not clear that an equilibrium model is appropriate, even for the central core of Sgr. For example Wang et al. (2022b) undertook a detailed study addressing this question using axisymmetric Jeans models, and showed that the core of Sgr may be reasonably well-described as in equilibrium. On the other hand, using DR2 data, Vasiliev & Belokurov (2020) find that the core is likely not in equilibrium, and therefore as an extension likely not well-described by Jeans-based dynamical models.
@@ 2.1 Coordinate transformation
$\begin{split}\cos\Gamma&=[\sin\delta\cos\delta_{0}\cos(\alpha-\alpha_{0})]/\sin\rho\\ \sin\Gamma&=[\cos\delta_{0}\sin(\alpha-\alpha_{0})]/\sin\rho.\end{split}$ (3)
@@ 2.2 Jeans Analysis
We utilize the spherical Jeans equation to connect the observed velocity dispersions to the underlying gravitational potential, ${\bf\Phi}$. We define $r$ as the 3-dimensional radial coordinate, and $R$ as the 2-dimensional projected radial coordinate. Note that $R$ is the magnitude of the vector ${\bf R}$ introduced above. In spherical coordinates the Jeans equation can be written as follows (Binney & Tremaine, 2008)
@@
where $\nu(r)$ is the stellar density profile, and the components of the intrinsic velocity dispersion are $\overline{u^{2}_{r}}(r),\overline{u^{2}_{\theta}}(r),\overline{u^{2}_{\phi}}(r)$. The stellar velocity anisotropy, $\beta_{a}$, is
@@
where $R_{HL}$ is the projected half light radius. To account for the contribution of the stellar mass to the gravitational potential, we assumes a stellar mass-to-light ratio of 1, and take the total luminosity of the Sgr core to be $L_{tot}=10^{7}L_{\odot}$, which is consistent with Majewski et al. (2003).
@@
This quantity is defined such that for an NFW profile, Equation 14 $\rightarrow 1$ in the center of the halo as $r\rightarrow 0$. On the other hand for a cored Burkert profile, this quantity $\rightarrow 0$ in the center of the halo as $r\rightarrow 0$. The log slope is useful because the shape parameters $\{M,c,\alpha,\beta,\gamma\}$ defining the dark matter profile are often degenerate, and are difficult to constrain from line-of-sight velocity data alone. It is possible to constrain more physical parameters such as mass enclosed within the half-light radius $M_{HL}$ and the log slope of the density profile at half-light radius (Strigari et al., 2007b; Guerra et al., 2023).
@@ 2.3 Bayesian analysis
For each of the velocity components, $\alpha=los,rad,ran$, we define an unbinned likelihood as
$\mathcal{L}_{\alpha}=\prod_{i=1}^{N}\frac{\exp\left[-\frac{1}{2}\frac{\left(u_{i,\alpha}-\langle u\rangle_{\alpha}\right)^{2}}{\delta_{\alpha,i}^{2}+\sigma^{2}_{\alpha}(R_{i})}\right]}{\sqrt{2\pi}\sqrt{\delta_{\alpha,i}^{2}+\sigma^{2}_{\alpha}(R_{i})}}.$ (15)
with the product being over the number of stars. The dispersion is the sum of the projected, intrinsic dispersion from the Jeans equation, and the measurement uncertainty of a given star, $\delta_{\alpha,\imath}$. The quantity $\langle u\rangle_{\alpha}$ represents the mean velocity of the stars in each component direction, after COM motion has been subtracted. This should be zero, or consistent with zero, in the case with no overall rotation or tidal stripping. For the case of the analysis of all three components, the likelihood is $L=\prod_{\alpha}L_{\alpha}$.
@@
The last important input is a value of projected half-light radius. Determination of the half-light radius for Sgr is complicated because the core blends into the tidal tails. Further, the estimate of the tidal radius depends on the stellar population considered. Previous estimates of the half-light radius range from $R_{HL}\simeq 2.6$ kpc in McConnachie (2012) and $R_{HL}\sim 1.1$ kpc in Ferguson & Strigari (2020). This latter value is also consistent with the determination from del Pino et al. (2021). For our analysis below, we use a half-light radius fit from our data that is consistent with the smaller of the two from previous literature.
@@ 3 Data
The left panel of Figure 1 shows a 2D histogram of all stars in proper motion space that pass the cuts described above. The Sgr region of proper motion space is identified by fitting a two component Gaussian mixture model to the selected stars in this region. Sgr member stars are identified as within the $2\sigma$ ellipse of the Gaussian Sgr component. The ellipse is given by means $\bm{\mu}\simeq(-2.695,-1.384)$[mas/yr], $\bm{\sigma}\simeq(0.129,0.160)$[mas/yr], semi major and minor axes=(0.640, 0.797)[mas/yr], and a counterclockwise rotation of $\theta=-0.861$ rad. For a more detailed discussion on similar methods for selecting member stars in dSphs, and the systematics associated with foreground subtraction, see Pace et al. (2022). Figure 1b shows the resulting $2\sigma$ ellipse that contains the Sgr component.
@@
To further minimize the number of field stars that may be confused with Sgr member stars, a line-of-sight velocity cut of $|v_{LOS}|<40$ km/s is applied after the coordinate transform described in Section 2.1, leaving 2729 candidate stars. Note that after the coordinate transformation, the mean velocity in each of the different component directions vanishes. We then fit a projected Plummer distribution to the resulting member star sample, and find the projected half light radius to be $2.6^{\circ}$, or $R_{HL}\approx 1.177$ kpc. This is consistent with estimates discussed above that have been obtained from Gaia data.
In a conservative effort to remove any possible effects of tidal disruption on our member sample, our final set of Sgr candidates includes only stars within this measured half-light radius of $2.6^{\circ}$. This results in 1404 Sgr candidates, which are shown in Figure 2. The 23 red points are those that are more than two times their error away from the expected mean parallax of Sgr, $\Pi_{Sgr}\sim 1/D_{0}$, and are removed from the candidate list. A color magnitude diagram of this final selection of 1381 stars is shown in Figure 3, and a list of Gaia DR3 source_ids for final selected candidates is published in the auxiliary file Sgr_candidates.txt.
@@
Measurements of parallaxes for stars at the distance of Sgr are neither precise nor accurate enough to utilize in our Jeans analysis below. Indeed, only one star in our sample has a parallax measured to less than 10%, i.e. $\delta\Pi/\Pi<0.1$. In light of this, to assign distances to the stars we generate two mock catalogs; one cataloge has an error on each distance “measurement” of $\delta_{D}=0.1$ kpc and one has $\delta_{D}=1$ kpc. To generate mock distances for each individual star, we sample from a normal distribution centered at the distance to the center of Sgr– 25.95 kpc (Majewski et al., 2003; del Pino et al., 2021)– and standard deviation of 1 kpc.
With these mean distances sampled for each star, to propagate errors, we sample values of $(\alpha,\delta,D,\mu_{\alpha*},\mu_{\delta},v_{\text{los}})$ 10,000 times from normal distributions centered at their measurement value with standard deviations of their measurement error. The coordinate transform is then performed for each of the 10,000 samples and the standard deviation of the resulting $(\rho,\phi_{\ast},v_{\text{los}},v_{\text{rad}},v_{\text{tan}})$ distributions is used as errors on those transformed parameters.
Figure 4 shows the resulting velocity dispersion for the different velocity components. Beyond the measured half-light radius, there is an increase in the line of sight velocity dispersion, the only component not influenced by mock distances. This inflation may be a result of contamination from stars in the tidal tails, or possibly a small number interloping stars that are not members of Sgr. This increase beyond the half-light radius justifies our sample cut which we use in the Jeans analysis below. One important thing to note here is that because the error bars shown in Figure 4 are jackknife estimated, not propagated errors, the $\delta_{D}=0.1$ kpc and $\delta_{D}=1$ kpc case are the same. However, changing the distribution of mock distances will change the velocity dispersion for $rad$ and $tan$ components. Figure 5 shows the effect of setting the distance of all stars to $D=D_{0}$, randomly selecting distances from a Gaussian distribution with standard deviation 0.5 kpc, and randomly selecting distances from a Gaussian distribution with standard deviation 1 kpc (as used in our analyzed sample).
@@ 4 Results
We perform separate MultiNest runs for the rad, tan, and los components for both the $\delta_{D}=0.1$ kpc and $\delta_{D}=1$ kpc datasets. For each dataset we also perform a run in which the likelihoods for the three velocity components are combined (a 6D analysis, abbreviated as Sgr All). As described above, in our runs we vary the shape parameters of the density profile, along with the scale mass, the scale concentration, and the velocity anisotropy, the latter of which is assumed to be constant over $R$. We then convert these quantities so that the reported posteriors are the parameters of velocity anisotropy $\log_{10}(1-\beta_{a})$, dark matter mass contained within the half light radius $\log_{10}(M_{HL}/M_{\odot})$, concentration $\log_{10}(c)$, and log slope at $R_{HL}$.
The resulting posteriors for each of the different runs are shown in Figure 7. In both the $\delta_{D}=0.1$ kpc and $\delta_{D}=1$ kpc cases, the 6D analyses show good agreement with LOS posteriors. As expected, there is a strong degeneracy between the log slope and the velocity anisotropy, which is particularly evident for the LOS-only data set. Also as expected, the mass within the half-light radius is well constrained by all of the data sets, and there is good agreement between the posteriors in all cases. The mass within the half-light radius is consistent between the two distance uncertainties. The measurement of the velocity anisotropy is also significantly improved with the inclusion of all three velocity components. We again emphasize that for the los data set the mean and the uncertainties on the posteriors reflect a measurement of the corresponding quantity. On the other hand, for the rad only, tan only, and all components runs, our analysis should be taken as a robust estimate of the expected measurement uncertainties, while the mean estimated quantities may be biased by the model for the distance uncertainties.
@@
Figure 7: Top corner plot shows posteriors for half light radius of 1.177 kpc and $\delta_{D}=0.1$ kpc. Bottom corner plot shows posteriors for half light radius of of 1.177 kpc and $\delta_{D}=1$ kpc. Black contours correspond to a 6D (all components) analysis, blue contours correspond to a line of sight velocity analysis only, orange contours correspond to radial velocities only, and green contours correspond to tangential velocities only. In both corner plots, the parameters are velocity anisotropy in the form $\log_{10}(1-\beta_{a})$, mass contained within the half light radius $\log_{10}(M_{HL}/M_{\odot})$, concentration $\log_{10}(c)$, and log slope at the half light radius LogSlope($r_{HL}$).
@@
As indicated in Figure 7, there are relatively weak constraints on the log-slope for an analysis which uses only any one of the three individual velocity components. However, there is significant improvement when using a likelihood that combines all three velocity components. In particular, in the analysis of the combined three data sets, the degeneracy between the log slope and the velocity anisotropy is broken, and a mild anisotropy is identified. In the LOS only case, which uses only real data, anisotropy is measured to be $\beta_{a}=-2.24\pm 1.99$, implying a tangentially biased system with large uncertainty. The log slope at $R_{HL}$ is $1.62\pm 0.69$. These errors improve in the full 6D analysis; however it is important to note that since mock distance values are used, this is an estimate of the constraining power of a 6D analysis rather than a measurement. For the $\delta_{D}=0.1$ kpc dataset, anisotropy is measured to be $\beta_{a}=-0.73\pm 0.18$ and log slope at $R_{HL}$ is $0.85\pm 0.13$. For the $\delta_{D}=1$ kpc dataset, anisotropy is measured to be $\beta_{a}=-0.28\pm 0.14$ and log slope at $R_{HL}$ is $0.74\pm 0.31$.
@@
Figure 8: Log slope as a function of spherical coordinates radius calculated from posteriors, for $\delta_{D}=0.1$ kpc (left column) and $\delta_{D}=1$ kpc (right column). Colored bands show the range of $1\sigma$ (top row) and $2\sigma$ (bottom row) value of log slope at a given $r$. Black with diagonal hatching represents the 6D (all components) Jeans analysis, blue with vertical hatching represents a line of sight velocity only analysis, orange with circle hatching represents a radial velocity only analysis, and green with star hatching represents a tangential velocity only analysis. The purple vertical dotted line is at the projected half light radius $R_{HL}=1.177$ kpc.

### v2

@@ Abstract
pubyear: 2025pagerange: Kinematics of the Sagittarius Dwarf Spheroidal core: A 5D Analysis for a 6D Methodology with Gaia DR3B
@@ 1 Introduction
The Sagittarius (Sgr) dwarf spheroidal (dSph) is one of the most luminous satellite galaxies orbiting the Milky Way (MW) (Ibata et al., 1994). It is unique from other dSphs in that it has both a central core and long, extended tails that indicate tidal interaction with the MW (Law & Majewski, 2010). The photometric and kinematic properties of the central core and tidal tails, combined with simulations of its orbit, provide insight into the nature of the Sgr progenitor before it was accreted into the MW, and provide estimates for its dark-to-luminous mass ratio (Peñarrubia et al., 2010; Łokas et al., 2010). Sgr-like progenitors have also been identified in large-volume cosmological simulations (Łokas, 2024), which provide insight into its evolution within a cosmological context.
@@
Though the dark matter distributions of many dSphs have been examined in detail, there have been relatively few studies of Sgr, particularly in the context of equilibrium models (Łokas et al., 2010). This is mainly because it is not clear that an equilibrium model is appropriate, even for the central core of Sgr. For example Wang et al. (2022b) undertook a detailed study addressing this question using axisymmetric Jeans models, and showed that the core of Sgr may be reasonably well-described as in equilibrium. On the other hand, using DR2 data, Vasiliev & Belokurov (2020) find that the core is likely not in equilibrium, and therefore as an extension likely not well-described by Jeans-based dynamical models.
@@ 2.1 Coordinate transformation
$\begin{split}\cos\Gamma&=[\sin\delta\cos\delta_{0}\cos(\alpha-\alpha_{0})-\cos{\delta}\sin{\delta_{0}}]/\sin\rho\\ \sin\Gamma&=[\cos\delta_{0}\sin(\alpha-\alpha_{0})]/\sin\rho.\end{split}$ (3)
@@ 2.2 Jeans Analysis
We utilize the spherical Jeans equation to connect the observed velocity dispersions to the underlying gravitational potential, ${\bf\Phi}$. We define $r$ as the 3-dimensional radial coordinate, and $R$ as the 2-dimensional projected radial coordinate. Note that $R$ is the magnitude of the vector ${\bf R}$ introduced above. In spherical coordinates and under the assumption of a time-independent spherical system the Jeans equation can be written as follows (Binney & Tremaine, 2008)
@@
where $\nu(r)$ is the stellar density profile, and the components of the intrinsic velocity dispersion are $\overline{u^{2}_{r}}(r),\overline{u^{2}_{\theta}}(r),\overline{u^{2}_{\phi}}(r)$. Note that the assumptions of a time-independent and spherical system mean we are not modeling streaming or tidal disruption. This also disregards the observed mean ellipticity $\epsilon\sim 0.64$ measured for Sgr (McConnachie, 2012). The stellar velocity anisotropy, $\beta_{a}$, is
@@
where $R_{HL}$ is the projected half light radius. To account for the contribution of the stellar mass to the gravitational potential, we assume a stellar mass-to-light ratio of 1. The actual stellar mass-to-light ratio of Sagittarius is not well constrained, but expectations for a typical dwarf galaxy and previous constraints place it around $\mathcal{O}(1)$ (Helmi & White, 2001; Vasiliev & Belokurov, 2020). We also take the total luminosity of the Sgr core to be $L_{tot}=10^{7}L_{\odot}$, which is consistent with Majewski et al. (2003).
@@
This quantity is defined such that for an NFW profile, Equation 14 $\rightarrow 1$ in the center of the halo as $r\rightarrow 0$. On the other hand for a cored Burkert profile, this quantity $\rightarrow 0$ in the center of the halo as $r\rightarrow 0$. The log slope is useful because the shape parameters $\{M,c,\alpha,\beta,\gamma\}$ defining the dark matter profile are often degenerate, and are difficult to constrain from line-of-sight velocity data alone (Binney & Mamon, 1982). It is possible to constrain more physical parameters such as mass enclosed within the half-light radius $M_{HL}$ and the log slope of the density profile at half-light radius (Strigari et al., 2007b; Guerra et al., 2023).
@@ 2.3 Bayesian analysis
For each of the velocity components, $comp=los,rad,ran$, we define an unbinned likelihood as
$\mathcal{L}_{comp}=\prod_{i=1}^{N}\frac{\exp\left[-\frac{1}{2}\frac{\left(u_{i,comp}-\langle u\rangle_{comp}\right)^{2}}{\Delta_{comp,i}^{2}+\sigma^{2}_{comp}(R_{i})}\right]}{\sqrt{2\pi}\sqrt{\Delta_{comp,i}^{2}+\sigma^{2}_{comp}(R_{i})}}.$ (15)
with the product being over the number of stars. The dispersion is the sum of the projected, intrinsic dispersion from the Jeans equation, and the measurement uncertainty of a given star, $\Delta_{comp,\imath}$. The quantity $\langle u\rangle_{comp}$ represents the mean velocity of the stars in each component direction, after COM motion has been subtracted. This should be zero, or consistent with zero, in the case with no overall rotation or tidal stripping as assumed in Section 2.2. A departure from zero would indicate that this assumption is not valid, and would not hold at larger radii where tidal stripping becomes important. It is for this reason that we have restricted our data to small radii. For the case of the analysis of all three components, the likelihood is $L=\prod_{comp}L_{comp}$.
@@
We tested larger prior ranges, including an increased $M_{200}$ range and allowing for negative $\gamma$, but saw no change in the resulting posteriors. The last important input is a value of projected half-light radius. Determination of the half-light radius for Sgr is complicated because the core blends into the tidal tails. Further, the estimate of the tidal radius depends on the stellar population considered. Previous estimates of the half-light radius range from $R_{HL}\simeq 2.6$ kpc in McConnachie (2012) and $R_{HL}\sim 1.1$ kpc in Ferguson & Strigari (2020). This latter value is also consistent with the determination from del Pino et al. (2021). For our analysis below, we use a half-light radius fit from our data that is consistent with the smaller of the two from previous literature.
@@ 3 Data
Figure 1 shows a 2D histogram of all stars in proper motion space returned by the original query with logarithmic normalization. The bottom left panel of Figure 2 shows a scatter plot in proper motion space of stars that pass the cuts described above. The Sgr region of proper motion space is identified by fitting a two component Gaussian mixture model to the selected stars in this region. Sgr member stars are identified as within the $2\sigma$ ellipse of the Gaussian Sgr component. The ellipse is given by means $\bm{\mu}\simeq(-2.695,-1.384)$[mas/yr], $\bm{\sigma}\simeq(0.129,0.160)$[mas/yr], semi major and minor axes=(0.640, 0.797)[mas/yr], and a counterclockwise rotation of $\theta=-0.861$ rad. For a more detailed discussion on similar methods for selecting member stars in dSphs, and the systematics associated with foreground subtraction, see Pace et al. (2022). The bottom left panel of Figure 2 shows the resulting $2\sigma$ ellipse that contains the Sgr component. The top left and bottom right panels show the projected PDFs plotted over a normalized histogram of the stars in $\mu_{\alpha*},\mu_{\delta}$ respectively, with lines corresponding to the Sgr and field components.
Figure 1: Proper motion space histogram of all stars returned by the Appendix A query, with logarithmic normalization.
@@
[…]
In a conservative effort to remove any possible effects of tidal disruption on our member sample, our final set of Sgr candidates includes only stars within this measured half-light radius of $2.6^{\circ}$. This results in 1404 Sgr candidates, which are shown in Figure 3. The 23 red points are those that are more than two times their error away from the expected mean parallax of Sgr, $\Pi_{Sgr}\sim 1/D_{0}$, and are removed from the candidate list. The minimum membership probability of stars within this final sample is $\sim 0.85$ as calculated by the Gaussian mixture model. A color magnitude diagram of this final selection of 1381 stars is shown in Figure 4, and a list of Gaia DR3 source_ids, with their associated membership probabilities as calculated by the Gaussian mixture model, for final selected candidates is published in the auxiliary file Sgr_candidates.txt.
@@
Measurements of parallaxes for stars at the distance of Sgr are neither precise nor accurate enough to utilize in our Jeans analysis below. Indeed, only one star in our sample has a parallax measured to less than 10%, i.e. $\Delta\Pi/\Pi<0.1$. In light of this, to assign distances to the stars we generate two mock catalogs; one catalog has an error on each distance “measurement” of $\Delta_{D}=0.1$ kpc and one has $\Delta_{D}=1$ kpc. To generate mock distances for each individual star, we sample from a normal distribution centered at the distance to the center of Sgr– 25.95 kpc (McConnachie, 2012; del Pino et al., 2021)– and standard deviation of 1 kpc.
With these mean distances sampled for each star, to propagate errors, we sample values of $(\alpha,\delta,D,\mu_{\alpha*},\mu_{\delta},v_{\text{los}})$ 10,000 times from normal distributions centered at their measurement value with standard deviations of their measurement error. The coordinate transform is then performed for each of the 10,000 samples and the standard deviation of the resulting $(\rho,\phi_{\ast},v_{\text{los}},v_{\text{rad}},v_{\text{tan}})$ distributions is used as errors on those transformed parameters. For simplicity we have not included the $\alpha,\delta,\mu_{\alpha*},\mu_{\delta}$ systematic errors and assumed them to be uncorrelated although both are present in Gaia measurements (Lindegren et al., 2021). No noticeable difference is found in RAD+TAN posteriors after applying the Gaia systematics correction for correlated proper motions found in Vasiliev & Baumgardt (2021). The details of that analysis are given in Appendix B.
Figure 5 shows the resulting velocity dispersion for the different velocity components. Beyond the measured half-light radius, there is an increase in the line of sight velocity dispersion, the only component not influenced by mock distances. This inflation may be a result of contamination from stars in the tidal tails, or possibly a small number interloping stars that are not members of Sgr. This increase beyond the half-light radius justifies our sample cut which we use in the Jeans analysis below. Note that the error bars shown in Figure 5 are jackknife resampling estimates made from the stars’ $v_{comp}$ value in a given bin, not propagated from the errors on those velocities, and as such will be the same for both $\Delta_{D}=0.1$ kpc and $\Delta_{D}=1$ kpc cases. However, changing the distribution of mock distances will change the velocity dispersion for $rad$ and $tan$ components. Figure 6 shows the effect of setting the distance of all stars to $D=D_{0}$, randomly selecting distances from a Gaussian distribution with standard deviation 0.5 kpc, and randomly selecting distances from a Gaussian distribution with standard deviation 1 kpc (as used in our analyzed sample).
@@ 4 Results
We perform separate MultiNest runs for the rad+tan combined (proper motion components) and los component for both the $\Delta_{D}=0.1$ kpc and $\Delta_{D}=1$ kpc datasets. For each dataset we also perform a run in which the likelihoods for all three velocity components are combined (a 6D analysis, abbreviated as Sgr All). As described above, in our runs we vary the shape parameters of the density profile, along with the scale mass, the scale concentration, and the velocity anisotropy, the latter of which is assumed to be constant over $R$. We then convert these quantities so that the reported posteriors are the parameters of velocity anisotropy $\log_{10}(1-\beta_{a})$, dark matter mass contained within the half light radius $\log_{10}(M_{HL}/M_{\odot})$, concentration $\log_{10}(c)$, and log slope at $R_{HL}$.
@@
Figure 8: Top corner plot shows posteriors for half light radius of 1.18 kpc and $\Delta_{D}=0.1$ kpc. Bottom corner plot shows posteriors for half light radius of of 1.18 kpc and $\Delta_{D}=1$ kpc. Black contours correspond to a 6D (all components) analysis, blue contours correspond to a line of sight velocity analysis only, and red contours correspond to a combined radial and tangential velocity analysis. In both corner plots, the parameters are velocity anisotropy in the form $\log_{10}(1-\beta_{a})$, mass contained within the half light radius $\log_{10}(M_{HL}/M_{\odot})$, concentration $\log_{10}(c)$, and log slope at the half light radius LogSlope($r_{HL}$).
The resulting posteriors for each of the different runs are shown in Figure 8. In both the $\Delta_{D}=0.1$ kpc and $\Delta_{D}=1$ kpc cases, the 6D analyses show good agreement with LOS posteriors. As expected, there is a strong degeneracy between the log slope and the velocity anisotropy, which is particularly evident for the LOS-only data set. Also as expected, the mass within the half-light radius is well constrained by all of the data sets, and there is good agreement between the posteriors in all cases. The mass within the half-light radius is consistent between the two distance uncertainties. The measurement of the velocity anisotropy is also significantly improved with the inclusion of all three velocity components. We again emphasize that for the los data set the mean and the uncertainties on the posteriors reflect a measurement of the corresponding quantity. On the other hand, for the rad and tan combined run, our analysis should be taken as a robust estimate of the expected measurement uncertainties, while the mean estimated quantities may be biased by the model for the distance uncertainties.
@@
As indicated in Figure 8, there are relatively weak constraints on the log-slope for an analysis which uses only any one of the three individual velocity components. However, there is significant improvement when using a likelihood that combines all three velocity components. In particular, in the analysis of the combined three data sets, the degeneracy between the log slope and the velocity anisotropy is broken, and a mild anisotropy is identified. This can also be seen by comparing the rad and tan run and the los only run posteriors. The addition of proper motions breaks the log slope and velocity anisotropy degeneracy. In the LOS only case, which uses only real data, anisotropy is measured to be $\beta_{a}=-2.24\pm 1.99$, implying a tangentially biased system with large uncertainty. This is consistent with previous measurements of a tangentially biased Sgr (Majewski et al., 2013). The log slope at $R_{HL}$ is $1.62\pm 0.69$. These errors improve in the full 6D analysis; however it is important to note that since mock distance values are used, this is an estimate of the constraining power of a 6D analysis rather than a measurement. For the $\Delta_{D}=0.1$ kpc dataset, anisotropy is measured to be $\beta_{a}=-0.73\pm 0.18$ and log slope at $R_{HL}$ is $0.85\pm 0.13$. For the $\Delta_{D}=1$ kpc dataset, anisotropy is measured to be $\beta_{a}=-0.28\pm 0.14$ and log slope at $R_{HL}$ is $0.74\pm 0.31$.
@@
Figure 9: Log slope as a function of spherical coordinates radius calculated from posteriors, for $\Delta_{D}=0.1$ kpc (left column) and $\Delta_{D}=1$ kpc (right column). Colored bands show the range of $1\sigma$ (top row) and $2\sigma$ (bottom row) value of log slope at a given $r$. Black with diagonal hatching represents the 6D (all components) Jeans analysis, blue with vertical hatching represents a line of sight velocity only analysis, orange with circle hatching represents a radial velocity only analysis, and green with star hatching represents a tangential velocity only analysis. The purple vertical dotted line is at the projected half light radius $R_{HL}=1.18$ kpc.